A creamery shop sells its special ice cream for $4.50 a quart. It costs them $3.00 a quart to make it. The daily demand for this flavor is normally distributed with a mean of 35 quarts and a standard deviation of 4 quarts. Unsold ice cream is sold each day to a local restaurant at $1.50 per quart. What is the service level and corresponding optimal stocking level?

95% service level; 50 quarts stock level

90% service level; 40 quarts stock level

65% service level; 30 quarts stock level

50% service level; 35 quarts stock level
Answer #1

1. D. 50% and 35 quarts

Explanation:

Cost of shortage(Cs) = sales price - cost price = 4.5 - 3 = 1.5

Cost of overage(Co) = Cost price - salvage value = 3 - 1.5 = 1.5

Service level = Cs / (Cs + Co) = 1.5 / (1.5 + 1.5) = 0.5

Z level that corresponds to 0.5 service level is 0

Optimal stocking level = Mean demand + Z * Standard deviation

= 35 + 0 *4 = 35